The difference of two integers is 9. Five times the smaller is 7 more than three times the larger. Find the numbers.
Answer by piglet162431(90)

Let n be the smaller number.
let n+9 be the bigger number.
5n=7+3(n+9)
5n=7+3n+27
5n=3n+34
n=17
Now u know the smaller number find the bigger number=26
Check
5(17)=7+3(26)
85=7+78
85=85
The 2 numbers are 26 and 17.